Identify the vertex of the parabola
y=4x^2+8x+6

Answer by stop12345now(6)

x coordinate of vertex is -(b/2a)= -(8/(4*2))=-8/8=-1
y coor- plug in x coor. into equation for x.
4(-1)^2+8(-1)+6
= 4-8+6
= 2
vertex is at (-1,2)